OpenStudy (anonymous):

In a 30-60-90 triangle, the shorter leg has length of 8√3 m. Find the length of the other leg (L) and the hypotenuse (H).

OpenStudy (anonymous):

ratios for this triangle are \(1:\sqrt{3}:2\)

Parth (parthkohli):

\(\Large \color{midnightblue}{\rightarrow x, x\sqrt3, 2x }\) 2x is the hypotenuse. xsqrt3 and x are two legs.

OpenStudy (anonymous):

so if short leg is \(8\sqrt{3}\) then long leg is \[8\sqrt{3}\times \sqrt{3}=8\times 3=24\]

Parth (parthkohli):

We know that xsqrt3 = 8sqrt3, so x = 8 Now find the values of 2x and x.